OpenStudy (anonymous):

limx→2 (x^2 − 5x + 6 / x^2 − 12x + 20 )= ? A) 1 B) ½ C) ¼ D) 1/8 E) 1/16

OpenStudy (anonymous):

\[\lim_{x\to2}\frac{x^2-5x+6}{x^2-12x+20}\\ \lim_{x\to2}\frac{(x-3)(x-2)}{(x-10)(x-2)}\\ \lim_{x\to2}\frac{x-3}{x-10}\]